OpenStudy (anonymous):

simplify: (3x + y) (2x - 1) - (3x + y)

OpenStudy (ageta):

Let's simplify step-by-step. (3x+y)(2x−1)−(3x+y) Distribute the Negative Sign: =(3x+y)(2x−1)−(3x+y) =(3x+y)(2x−1)+−3x+−y =(3x+y)(2x−1)+−3x+−y Distribute: =(3x)(2x)+(3x)(−1)+(y)(2x)+(y)(−1)+−3x+−y =6x2+−3x+2xy+−y+−3x+−y Combine Like Terms: =6x2+−3x+2xy+−y+−3x+−y =(6x2)+(2xy)+(−3x+−3x)+(−y+−y) =6x2+2xy+−6x+−2y Answer: =6x2+2xy−6x−2y

OpenStudy (12man):

Let's simplify step-by-step. (3x+y)(2x−1)−(3x+y) Distribute the Negative Sign: =(3x+y)(2x−1)−(3x+y) =(3x+y)(2x−1)+−3x+−y =(3x+y)(2x−1)+−3x+−y Distribute: =(3x)(2x)+(3x)(−1)+(y)(2x)+(y)(−1)+−3x+−y =6x2+−3x+2xy+−y+−3x+−y Combine Like Terms: =6x2+−3x+2xy+−y+−3x+−y =(6x2)+(2xy)+(−3x+−3x)+(−y+−y) =6x2+2xy+−6x+−2y Answer: =6x2+2xy−6x−2y

OpenStudy (ageta):

you just copy mine

OpenStudy (anonymous):

kk :)

OpenStudy (ageta):

can i get a point

OpenStudy (ageta):

tnxs